# Upper Approximation Bounds for Neural Oscillators

### v1

@@ B.1 Proof of Lemma 5
$\begin{array}[]{l}{{\bf{z}}_{2}}(t,{{\bf{z}}_{0}})-{{\bf{z}}_{1}}(t,{{\bf{z}}_{0}})\\ \;\;=\int_{0}^{t}{\left\{{\int_{\tau}^{t}{\frac{{\partial{{{\bf{z}}}_{1}^{\prime}}\left[{\lambda,{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}{{\partial{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}}{\rm{d}}\lambda}+{\bf{I}}}\right\}}{{\begin{Bmatrix}{\bf{0}}\\ {{g_{2}}\left[{{{\bf{z}}_{2,1}}(\tau,{{\bf{z}}_{0}}),{{\bf{z}}_{2,2}}(\tau,{{\bf{z}}_{0}}),{\bf{u}}(\tau)}\right]-{g_{1}}\left[{{{\bf{z}}_{2,1}}(\tau,{{\bf{z}}_{0}}),{{\bf{z}}_{2,2}}(\tau,{{\bf{z}}_{0}}),{\bf{u}}(\tau)}\right]}\end{Bmatrix}}}\rm{d}\tau\\ \;\;\mathop{=}\limits^{(a)}\int_{0}^{t}{\frac{{\partial{{\bf{z}}_{1}}\left[{t,{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}{{\partial{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}}{\begin{Bmatrix}{\bf{0}}\\ {{g_{2}}\left[{{{\bf{z}}_{2,1}}(\tau,{{\bf{z}}_{0}}),{{\bf{z}}_{2,2}}(\tau,{{\bf{z}}_{0}}),{\bf{u}}(\tau)}\right]-{g_{1}}\left[{{{\bf{z}}_{2,1}}(\tau,{{\bf{z}}_{0}}),{{\bf{z}}_{2,2}}(\tau,{{\bf{z}}_{0}}),{\bf{u}}(\tau)}\right]}\end{Bmatrix}}{\rm{d}}\tau}\\ \;\;=\int_{0}^{t}{\frac{{\partial{{\bf{z}}_{1}}\left[{t,{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}{{\partial{{\bf{z}}_{2,2}}(\tau,{{\bf{z}}_{0}})}}\left\{{{g_{2}}\left[{{{\bf{z}}_{2,1}}(\tau,{{\bf{z}}_{0}}),{{\bf{z}}_{2,2}}(\tau,{{\bf{z}}_{0}}),{\bf{u}}(\tau)}\right]-{g_{1}}\left[{{{\bf{z}}_{2,1}}(\tau,{{\bf{z}}_{0}}),{{\bf{z}}_{2,2}}(\tau,{{\bf{z}}_{0}}),{\bf{u}}(\tau)}\right]}\right\}{\rm{d}}\tau},\end{array}$ (83)
@@ B.2 Proof of Lemma 6
[…]
@@
$\begin{array}[]{l}\left|{{{\bf{z}}_{2}}(t,{{\bf{z}}_{0}})-{{\bf{z}}_{1}}(t,{{\bf{z}}_{0}})}\right|\\ \;\;\leq\int_{0}^{t}{\left|{\exp\left\{{\int_{\tau}^{t}{\frac{{\partial{{{\bf{z}}}_{1}^{\prime}}\left[{\hat{\tau},{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}{{\partial{{\bf{z}}_{1}}\left[{\hat{\tau},{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}{\rm{d}}\hat{\tau}}}\right\}{\bf{\Delta}}g(\tau,{{\bf{z}}_{0}})}\right|{\rm{d}}\tau}=\int_{0}^{t}{\left|{\mathop{\lim}\limits_{\scriptstyle\Delta\hat{\tau}\to 0\atop\scriptstyle n\to+\infty}\prod\limits_{i=1}^{n}{\exp\left\{{\frac{{\partial{{{\bf{z}}}_{1}^{\prime}}\left[{{{\hat{\tau}}_{i}},{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}{{\partial{{\bf{z}}_{1}}\left[{{{\hat{\tau}}_{i}},{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}\Delta\hat{\tau}}\right\}}{\bf{\Delta}}g(\tau,{{\bf{z}}_{0}})}\right|{\rm{d}}\tau}\\ \;\;\leq\int_{0}^{t}{\mathop{\lim}\limits_{\scriptstyle\Delta\hat{\tau}\to 0\atop\scriptstyle n\to+\infty}\left|{\prod\limits_{i=1}^{n}{\left({\sum\limits_{m=0}^{+\infty}{\frac{1}{{m!}}}{{\left\{{\frac{{\partial{{{\bf{z}}}_{1}^{\prime}}\left[{{{\hat{\tau}}_{i}},{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}{{\partial{{\bf{z}}_{1}}\left[{{{\hat{\tau}}_{i}},{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}\Delta\hat{\tau}}\right\}}^{m}}}\right)}{\bf{\Delta}}g(\tau,{{\bf{z}}_{0}})}\right|{\rm{d}}\tau}\\ \;\;=\int_{0}^{t}{\mathop{\lim}\limits_{\scriptstyle\Delta\hat{\tau}\to 0\atop\scriptstyle n\to+\infty}\left|{\left({\sum\limits_{m=0}^{+\infty}{\frac{1}{{m!}}}{{\left\{{\frac{{\partial{{{\bf{z}}}_{1}^{\prime}}\left[{{{\hat{\tau}}_{n}},{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}{{\partial{{\bf{z}}_{1}}\left[{{{\hat{\tau}}_{n}},{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}\Delta\hat{\tau}}\right\}}^{m}}}\right)\prod\limits_{i=1}^{n-1}{\left({\sum\limits_{m=0}^{+\infty}{\frac{1}{{m!}}}{{\left\{{\frac{{\partial{{{\bf{z}}}_{1}^{\prime}}\left[{{{\hat{\tau}}_{i}},{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}{{\partial{{\bf{z}}_{1}}\left[{{{\hat{\tau}}_{i}},{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}\Delta\hat{\tau}}\right\}}^{m}}}\right)}{\bf{\Delta}}g(\tau,{{\bf{z}}_{0}})}\right|{\rm{d}}\tau}\\ \;\;\mathop{\leq}\limits^{(a)}\int_{0}^{t}{\mathop{\lim}\limits_{\scriptstyle\Delta\hat{\tau}\to 0\atop\scriptstyle n\to+\infty}\left[{\sum\limits_{m=0}^{+\infty}{\frac{1}{{m!}}}{{\left({{L_{{g_{1}}}}\Delta\hat{\tau}}\right)}^{m}}}\right]\left|{\prod\limits_{i=1}^{n-1}{\left({\sum\limits_{m=0}^{+\infty}{\frac{1}{{m!}}}{{\left\{{\frac{{\partial{{{\bf{z}}}_{1}^{\prime}}\left[{{{\hat{\tau}}_{i}},{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}{{\partial{{\bf{z}}_{1}}\left[{{{\hat{\tau}}_{i}},{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}\Delta\hat{\tau}}\right\}}^{m}}}\right)}{\bf{\Delta}}g(\tau,{{\bf{z}}_{0}})}\right|{\rm{d}}\tau}\\ \;\;\leq\int_{0}^{t}{\mathop{\lim}\limits_{\scriptstyle\Delta\hat{\tau}\to 0\atop\scriptstyle n\to+\infty}{{\left[{\sum\limits_{m=0}^{+\infty}{\frac{1}{{m!}}}{{\left({{L_{{g_{1}}}}\Delta\hat{\tau}}\right)}^{m}}}\right]}^{n}}\left|{{\bf{\Delta}}g(\tau,{{\bf{z}}_{0}})}\right|{\rm{d}}\tau}=\int_{0}^{t}{\exp\left({\int_{\tau}^{t}{{L_{{g_{1}}}}{\rm{d}}\hat{\tau}}}\right)\left|{{\bf{\Delta}}g(\tau,{{\bf{z}}_{0}})}\right|{\rm{d}}\tau}\\ \;\;\leq{e^{T{L_{{g_{1}}}}}}T\mathop{\max}\limits_{\tau\in[0,T]}\left|{{\bf{\Delta}}g(\tau,{{\bf{z}}_{0}})}\right|,\end{array}$ (94)
@@
$\begin{array}[]{l}\left|{{{\bf{z}}_{2}}(t,{{\bf{z}}_{0}})-{{\bf{z}}_{1}}(t,{{\bf{z}}_{0}})}\right|\\ \;\;\leq\int_{0}^{t}{\left|{\frac{{\partial{{\bf{z}}_{1}}\left[{t,{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}{{\partial{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}}{\begin{Bmatrix}{\mathbf{0}}_{r\times 1}\\ {{g_{2}}\left[{{{\bf{z}}_{2,1}}(\tau,{{\bf{z}}_{0}}),{{\bf{z}}_{2,2}}(\tau,{{\bf{z}}_{0}}),{\bf{u}}(\tau)}\right]-{g_{1}}\left[{{{\bf{z}}_{2,1}}(\tau,{{\bf{z}}_{0}}),{{\bf{z}}_{2,2}}(\tau,{{\bf{z}}_{0}}),{\bf{u}}(\tau)}\right]}\end{Bmatrix}}}\right|{\rm{d}}\tau}\\ \;\;\leq\int_{0}^{t}{{\left|{\left\{{\frac{{\partial{{\bf{z}}_{1}}\left[{t,{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}{{\partial{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}}}\right\}^{\top}}\right|}_{2,1}{{\left|{{g_{2}}\left[{{{\bf{z}}_{2,1}}(\tau,{{\bf{z}}_{0}}),{{\bf{z}}_{2,2}}(\tau,{{\bf{z}}_{0}}),{\bf{u}}(\tau)}\right]-{g_{1}}\left[{{{\bf{z}}_{2,1}}(\tau,{{\bf{z}}_{0}}),{{\bf{z}}_{2,2}}(\tau,{{\bf{z}}_{0}}),{\bf{u}}(\tau)}\right]}\right|}_{2}}{\rm{d}}\tau}\\ \;\;\leq\mathop{\max}\limits_{\tau\in[0,T]}{\left|{{\bf{\Delta}}g(\tau,{{\bf{z}}_{0}})}\right|_{2}}\int_{0}^{t}{{{\left|\left\{{\frac{{\partial{{\bf{z}}_{1}}\left[{t,{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}{{\partial{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}}}\right\}^{\top}\right|}_{2,1}}{\rm{d}}\tau},\end{array}$ (95)
@@
$\begin{array}[]{l}{\left|\left\{{\frac{{\partial{{\bf{z}}_{1}}\left[{t,{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}{{\partial{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}}}\right\}^{\top}\right|_{2,1}}=\sum\limits_{k=1}^{2r}{{{\left|{\frac{{\partial{z_{1,k}}\left({t,{{\bf{z}}_{2}}}\right)}}{{\partial{{\bf{z}}_{2}}}}}\right|}_{2}}}=\sum\limits_{k=1}^{2r}{\left|{\frac{{\partial{z_{1,k}}\left({t,{{\bf{z}}_{2}}}\right)}}{{\partial{{\bf{z}}_{2}}}}\cdot{{\bf{\Delta}}_{k}}}\right|}\\ \;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;=\sum\limits_{k=1}^{2r}{\mathop{\lim}\limits_{v\to 0}\frac{{\left|{{z_{1,k}}\left({t,{{\bf{z}}_{2}}+v{{\bf{\Delta}}_{k}}}\right)-{z_{1,k}}\left({t,{{\bf{z}}_{2}}}\right)}\right|}}{{{{\left|{v{{\bf{\Delta}}_{k}}}\right|}_{2}}}}}\\ \;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\mathop{\leq}\limits^{(a)}\sum\limits_{k=1}^{2r}{\mathop{\lim}\limits_{v\to 0}\frac{{{\beta_{k}}({{\left|{v{{\bf{\Delta}}_{k}}}\right|}_{2}},t-\tau)}}{{{{\left|{v{{\bf{\Delta}}_{k}}}\right|}_{2}}}}}={\left.{\sum\limits_{k=1}^{2r}{\frac{\partial}{{\partial h}}{\beta_{k}}(h,t-\tau)}}\right|_{h={0^{+}}}},\end{array}$ (96)
@@
$\begin{array}[]{l}\left|{{{\bf{z}}_{2}}(t,{{\bf{z}}_{0}})-{{\bf{z}}_{1}}(t,{{\bf{z}}_{0}})}\right|\leq\mathop{\max}\limits_{\tau\in[0,T]}{\left|{{\bf{\Delta}}g(\tau,{{\bf{z}}_{0}})}\right|_{2}}\int_{0}^{t}{{\left.{\sum\limits_{k=1}^{2r}{\frac{\partial}{{\partial h}}{\beta_{k}}(h,t-\tau)}}\right|_{h={0^{+}}}}{\rm{d}}\tau}\\ \;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\mathop{\leq}\limits^{(a)}\mathop{\max}\limits_{\tau\in[0,T]}{\left|{{\bf{\Delta}}g(\tau,{{\bf{z}}_{0}})}\right|_{2}}{B_{\beta}},\end{array}$ (97)

### v2

@@ B.1 Proof of Lemma 5
${{\bf{z}}_{2}}(t,{{\bf{z}}_{0}})-{{\bf{z}}_{1}}(t,{{\bf{z}}_{0}})$ (88)
$\;\;=\int_{0}^{t}{\left\{{\int_{\tau}^{t}{\frac{{\partial{{{\bf{z}}}_{1}^{\prime}}\left[{\lambda,{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}{{\partial{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}}{\rm{d}}\lambda}+{\bf{I}}}\right\}}{{\begin{Bmatrix}{\bf{0}}\\ {{g_{2}}\left[{{{\bf{z}}_{2,1}}(\tau,{{\bf{z}}_{0}}),{{\bf{z}}_{2,2}}(\tau,{{\bf{z}}_{0}}),{\bf{u}}(\tau)}\right]-{g_{1}}\left[{{{\bf{z}}_{2,1}}(\tau,{{\bf{z}}_{0}}),{{\bf{z}}_{2,2}}(\tau,{{\bf{z}}_{0}}),{\bf{u}}(\tau)}\right]}\end{Bmatrix}}}\rm{d}\tau$
$\;\;\mathop{=}\limits^{(a)}\int_{0}^{t}{\frac{{\partial{{\bf{z}}_{1}}\left[{t,{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}{{\partial{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}}{\begin{Bmatrix}{\bf{0}}\\ {{g_{2}}\left[{{{\bf{z}}_{2,1}}(\tau,{{\bf{z}}_{0}}),{{\bf{z}}_{2,2}}(\tau,{{\bf{z}}_{0}}),{\bf{u}}(\tau)}\right]-{g_{1}}\left[{{{\bf{z}}_{2,1}}(\tau,{{\bf{z}}_{0}}),{{\bf{z}}_{2,2}}(\tau,{{\bf{z}}_{0}}),{\bf{u}}(\tau)}\right]}\end{Bmatrix}}{\rm{d}}\tau}$
$\;\;=\int_{0}^{t}{\frac{{\partial{{\bf{z}}_{1}}\left[{t,{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}{{\partial{{\bf{z}}_{2,2}}(\tau,{{\bf{z}}_{0}})}}\left\{{{g_{2}}\left[{{{\bf{z}}_{2,1}}(\tau,{{\bf{z}}_{0}}),{{\bf{z}}_{2,2}}(\tau,{{\bf{z}}_{0}}),{\bf{u}}(\tau)}\right]-{g_{1}}\left[{{{\bf{z}}_{2,1}}(\tau,{{\bf{z}}_{0}}),{{\bf{z}}_{2,2}}(\tau,{{\bf{z}}_{0}}),{\bf{u}}(\tau)}\right]}\right\}{\rm{d}}\tau},$
@@ B.2 Proof of Lemma 6
$\frac{{\partial{{{\bf{z}}}_{1}^{\prime}}\left[{t,{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}{{\partial{{\bf{z}}_{1}}\left[{t,{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}={\begin{bmatrix}{{\bf{0}}}_{r\times r}&{{{\bf{I}}_{r\times r}}}\\ {\frac{{\partial{g_{1}}\left\{{{{\bf{z}}_{1,1}}\left[{t,{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right],{{\bf{z}}_{1,2}}\left[{t,{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right],{\bf{u}}(t)}\right\}}}{{\partial{{\bf{z}}_{1,1}}\left[{t,{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}}&{\frac{{\partial{g_{1}}\left\{{{{\bf{z}}_{1,1}}\left[{t,{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right],{{\bf{z}}_{1,2}}\left[{t,{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right],{\bf{u}}(t)}\right\}}}{{\partial{{\bf{z}}_{1,2}}\left[{t,{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}}\end{bmatrix}}$. (90)
@@
$\;\;\leq\int_{0}^{t}{\left|{\exp\left\{{\int_{\tau}^{t}{\frac{{\partial{{{\bf{z}}}_{1}^{\prime}}\left[{\hat{\tau},{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}{{\partial{{\bf{z}}_{1}}\left[{\hat{\tau},{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}{\rm{d}}\hat{\tau}}}\right\}{\bf{\Delta}}g(\tau,{{\bf{z}}_{0}})}\right|{\rm{d}}\tau}=\int_{0}^{t}{\left|{\mathop{\lim}\limits_{\scriptstyle\Delta\hat{\tau}\to 0\atop\scriptstyle n\to+\infty}\prod\limits_{i=1}^{n}{\exp\left\{{\frac{{\partial{{{\bf{z}}}_{1}^{\prime}}\left[{{{\hat{\tau}}_{i}},{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}{{\partial{{\bf{z}}_{1}}\left[{{{\hat{\tau}}_{i}},{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}\Delta\hat{\tau}}\right\}}{\bf{\Delta}}g(\tau,{{\bf{z}}_{0}})}\right|{\rm{d}}\tau}$
$\;\;\leq\int_{0}^{t}{\mathop{\lim}\limits_{\scriptstyle\Delta\hat{\tau}\to 0\atop\scriptstyle n\to+\infty}\left|{\prod\limits_{i=1}^{n}{\left({\sum\limits_{m=0}^{+\infty}{\frac{1}{{m!}}}{{\left\{{\frac{{\partial{{{\bf{z}}}_{1}^{\prime}}\left[{{{\hat{\tau}}_{i}},{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}{{\partial{{\bf{z}}_{1}}\left[{{{\hat{\tau}}_{i}},{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}\Delta\hat{\tau}}\right\}}^{m}}}\right)}{\bf{\Delta}}g(\tau,{{\bf{z}}_{0}})}\right|{\rm{d}}\tau}$
$\;\;=\int_{0}^{t}{\mathop{\lim}\limits_{\scriptstyle\Delta\hat{\tau}\to 0\atop\scriptstyle n\to+\infty}\left|{\left({\sum\limits_{m=0}^{+\infty}{\frac{1}{{m!}}}{{\left\{{\frac{{\partial{{{\bf{z}}}_{1}^{\prime}}\left[{{{\hat{\tau}}_{n}},{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}{{\partial{{\bf{z}}_{1}}\left[{{{\hat{\tau}}_{n}},{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}\Delta\hat{\tau}}\right\}}^{m}}}\right)\prod\limits_{i=1}^{n-1}{\left({\sum\limits_{m=0}^{+\infty}{\frac{1}{{m!}}}{{\left\{{\frac{{\partial{{{\bf{z}}}_{1}^{\prime}}\left[{{{\hat{\tau}}_{i}},{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}{{\partial{{\bf{z}}_{1}}\left[{{{\hat{\tau}}_{i}},{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}\Delta\hat{\tau}}\right\}}^{m}}}\right)}{\bf{\Delta}}g(\tau,{{\bf{z}}_{0}})}\right|{\rm{d}}\tau}$
$\;\;\mathop{\leq}\limits^{(a)}\int_{0}^{t}{\mathop{\lim}\limits_{\scriptstyle\Delta\hat{\tau}\to 0\atop\scriptstyle n\to+\infty}\left[{\sum\limits_{m=0}^{+\infty}{\frac{1}{{m!}}}{{\left({{L_{{g_{1}}}}\Delta\hat{\tau}}\right)}^{m}}}\right]\left|{\prod\limits_{i=1}^{n-1}{\left({\sum\limits_{m=0}^{+\infty}{\frac{1}{{m!}}}{{\left\{{\frac{{\partial{{{\bf{z}}}_{1}^{\prime}}\left[{{{\hat{\tau}}_{i}},{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}{{\partial{{\bf{z}}_{1}}\left[{{{\hat{\tau}}_{i}},{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}\Delta\hat{\tau}}\right\}}^{m}}}\right)}{\bf{\Delta}}g(\tau,{{\bf{z}}_{0}})}\right|{\rm{d}}\tau}$
$\;\;\leq\int_{0}^{t}{\mathop{\lim}\limits_{\scriptstyle\Delta\hat{\tau}\to 0\atop\scriptstyle n\to+\infty}{{\left[{\sum\limits_{m=0}^{+\infty}{\frac{1}{{m!}}}{{\left({{L_{{g_{1}}}}\Delta\hat{\tau}}\right)}^{m}}}\right]}^{n}}\left|{{\bf{\Delta}}g(\tau,{{\bf{z}}_{0}})}\right|{\rm{d}}\tau}=\int_{0}^{t}{\exp\left({\int_{\tau}^{t}{{L_{{g_{1}}}}{\rm{d}}\hat{\tau}}}\right)\left|{{\bf{\Delta}}g(\tau,{{\bf{z}}_{0}})}\right|{\rm{d}}\tau}$
$\;\;\leq{e^{T{L_{{g_{1}}}}}}T\mathop{\max}\limits_{\tau\in[0,T]}\left|{{\bf{\Delta}}g(\tau,{{\bf{z}}_{0}})}\right|,$
@@
$\left|{{{\bf{z}}_{2}}(t,{{\bf{z}}_{0}})-{{\bf{z}}_{1}}(t,{{\bf{z}}_{0}})}\right|$ (100)
$\;\;\leq\int_{0}^{t}{\left|{\frac{{\partial{{\bf{z}}_{1}}\left[{t,{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}{{\partial{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}}{\begin{Bmatrix}{\mathbf{0}}_{r\times 1}\\ {{g_{2}}\left[{{{\bf{z}}_{2,1}}(\tau,{{\bf{z}}_{0}}),{{\bf{z}}_{2,2}}(\tau,{{\bf{z}}_{0}}),{\bf{u}}(\tau)}\right]-{g_{1}}\left[{{{\bf{z}}_{2,1}}(\tau,{{\bf{z}}_{0}}),{{\bf{z}}_{2,2}}(\tau,{{\bf{z}}_{0}}),{\bf{u}}(\tau)}\right]}\end{Bmatrix}}}\right|{\rm{d}}\tau}$
$\;\;\leq\int_{0}^{t}{{\left|{\left\{{\frac{{\partial{{\bf{z}}_{1}}\left[{t,{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}{{\partial{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}}}\right\}^{\top}}\right|}_{2,1}{{\left|{{g_{2}}\left[{{{\bf{z}}_{2,1}}(\tau,{{\bf{z}}_{0}}),{{\bf{z}}_{2,2}}(\tau,{{\bf{z}}_{0}}),{\bf{u}}(\tau)}\right]-{g_{1}}\left[{{{\bf{z}}_{2,1}}(\tau,{{\bf{z}}_{0}}),{{\bf{z}}_{2,2}}(\tau,{{\bf{z}}_{0}}),{\bf{u}}(\tau)}\right]}\right|}_{2}}{\rm{d}}\tau}$
$\;\;\leq\mathop{\max}\limits_{\tau\in[0,T]}{\left|{{\bf{\Delta}}g(\tau,{{\bf{z}}_{0}})}\right|_{2}}\int_{0}^{t}{{{\left|\left\{{\frac{{\partial{{\bf{z}}_{1}}\left[{t,{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}{{\partial{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}}}\right\}^{\top}\right|}_{2,1}}{\rm{d}}\tau},$
@@
${\left|\left\{{\frac{{\partial{{\bf{z}}_{1}}\left[{t,{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}{{\partial{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}}}\right\}^{\top}\right|_{2,1}}=\sum\limits_{k=1}^{2r}{{{\left|{\frac{{\partial{z_{1,k}}\left({t,{{\bf{z}}_{2}}}\right)}}{{\partial{{\bf{z}}_{2}}}}}\right|}_{2}}}=\sum\limits_{k=1}^{2r}{\left|{\frac{{\partial{z_{1,k}}\left({t,{{\bf{z}}_{2}}}\right)}}{{\partial{{\bf{z}}_{2}}}}\cdot{{\bf{\Delta}}_{k}}}\right|}$ (101)
$\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;=\sum\limits_{k=1}^{2r}{\mathop{\lim}\limits_{v\to 0}\frac{{\left|{{z_{1,k}}\left({t,{{\bf{z}}_{2}}+v{{\bf{\Delta}}_{k}}}\right)-{z_{1,k}}\left({t,{{\bf{z}}_{2}}}\right)}\right|}}{{{{\left|{v{{\bf{\Delta}}_{k}}}\right|}_{2}}}}}$
$\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\mathop{\leq}\limits^{(a)}\sum\limits_{k=1}^{2r}{\mathop{\lim}\limits_{v\to 0}\frac{{{\beta_{k}}({{\left|{v{{\bf{\Delta}}_{k}}}\right|}_{2}},t-\tau)}}{{{{\left|{v{{\bf{\Delta}}_{k}}}\right|}_{2}}}}}={\left.{\sum\limits_{k=1}^{2r}{\frac{\partial}{{\partial h}}{\beta_{k}}(h,t-\tau)}}\right|_{h={0^{+}}}},$
@@
$\left|{{{\bf{z}}_{2}}(t,{{\bf{z}}_{0}})-{{\bf{z}}_{1}}(t,{{\bf{z}}_{0}})}\right|\leq\mathop{\max}\limits_{\tau\in[0,T]}{\left|{{\bf{\Delta}}g(\tau,{{\bf{z}}_{0}})}\right|_{2}}\int_{0}^{t}{{\left.{\sum\limits_{k=1}^{2r}{\frac{\partial}{{\partial h}}{\beta_{k}}(h,t-\tau)}}\right|_{h={0^{+}}}}{\rm{d}}\tau}\mathop{\leq}\limits^{(a)}\mathop{\max}\limits_{\tau\in[0,T]}{\left|{{\bf{\Delta}}g(\tau,{{\bf{z}}_{0}})}\right|_{2}}{B_{\beta}},$ (102)